OpenStudy (anonymous):

Heidi left a 10% tip of $4.60 on a meal. What was the cost of the meal before the tip? What was the total cost of the meal including the tip?

OpenStudy (anonymous):

10 times 4.60?

OpenStudy (anonymous):

no, umm justa sec i needa think lol

OpenStudy (anonymous):

kk:)

OpenStudy (anonymous):

Ok its 450 *10

OpenStudy (anonymous):

ya u were right lol i had to reread it

OpenStudy (anonymous):

but thats for the total cost..how do i find the cost before the tip was added onto the orginal price?

OpenStudy (anonymous):

Before tip was added=$45.00 After tip was added=$49.50

OpenStudy (anonymous):

but how do i get the price before so i can understand it/

OpenStudy (anonymous):

the original price, how did you get the 45?

OpenStudy (anonymous):

Before tip was added= 10*45 After tip was added= 10*45+4.50

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

ok np